OpenStudy (anonymous):

How many Carbons and Hydrogens are in 3-methyl 4-propyl heptane?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

Methe = 1 carbon Prop= 3 carbons Hept= 7 carbons -ane - Alkanes so the General Formula is Cn H2n + 2 So, Carbon = 11 Hydrogen = 11 * 2 + 2 = 24